OpenStudy (anonymous):

A rocket sled is tested at "5 g" (5 times the acceleration due to gravity). If the sled starts from rest at position do= 0.00, how long does it take to travel 441 meters? t = _____ s

OpenStudy (anonymous):

d= 1/2*a*t^2 441 = 1/2*5g*t^2 solve for t